# Cendol: Open Instruction-tuned Generative Large Language Models for Indonesian Languages

### v1

@@ 1 Introduction
Indonesia is the fourth most populous country in the world, with around 280 million people spread across more than 17,000 islands within a humongous area of 2 million square kilometers. With such a large archipelago surrounding the country, digital services become immensely crucial making Indonesia the fourth largest internet user in the world, with 220 million users. Despite the huge demand, the technology supporting Indonesian digital businesses still lags compared to other much smaller countries. One aspect that it still left behind is the access to state-of-the-art large language model (LLM) technology, such as ChatGPT OpenAI (2023a) and GPT4 OpenAI (2023b). Although these LLMs support Indonesian and its local languages, these LLMs often have much weaker language representation for such low-resource and underrepresented languages Cahyawijaya et al. (2023b, a); Asai et al. (2023).
@@ 2.1 Indonesian Language Models
Various pre-trained Indonesian language models (LMs) have emerged in the past years, including IndoBERT Wilie et al. (2020); Koto et al. (2020, 2021), IndoBART Cahyawijaya et al. (2021), and IndoGPT Cahyawijaya et al. (2021). These models have a smaller parameter size compared to recent LLMs and have primarily been evaluated only on standard NLP benchmarks. Concurrently, advancements in LLMs have led to the development of multilingual LLMs like BLOOM Scao et al. (2022) and mT5 Xue et al. (2021), which include Indonesian, Javanese, and Sundanese. Yet, they fall short of covering other underrepresented Indonesian local languages. LLaMA-2 Touvron et al. (2023b) also incorporates Indonesian, although it comprises a small portion (0.03%), diminishing its usability in the Indonesian context. Additionally, multilingual LLMs focusing on Southeast Asian languages, such as SEA-LION Singapore (2023) and SeaLLM Nguyen et al. (2023), are beginning to rise, indicating an increasing demand for refining LLMs for underrepresented languages.
@@ 3.3 Local Language Generative Prompt
To cover more underrepresented languages spoken in Indonesia, we collect Indonesian local language prompts from two sources, i.e., Wikipedia in local languages and NusaParagraph Cahyawijaya et al. (2023b). We covered 18 local languages including Sundanese (sun), Javanese (jav), Acehnese (ace), Banjarese (bjn), Buginese (bug), and Gorontalo (gor). Since many local languages in Indonesia are derived from standard Malaysia (msa), we also collect the prompt from Malaysian Wikipedia.777https://ms.wikipedia.org For the prompt from Wikipedia, we incorporate the same prompt generation strategy as in §3.2, while for the generative prompt from NusaParagraph, we invert the input and output label of the dataset to make a sentence generation task for the specified local language. In total, we collect 6.27M local language generative prompts.
@@ 4.1 Backbone Models
We prepare Cendol models of various base models to enable thorough comparison and analysis across different scales. Specifically, we train Cendol from models of different sizes, from 300M up to 13B parameters, to see the impact of size on performance. We also explore using decoder-only and encoder-only models, and lastly, we also explore using models of different origins to see if the starting base model has any impact. Finally, by producing Cendol in different configurations, users can choose their models to use based on needs and constraints. Specifically, we train Cendol by continuously fine-tuning decoder-only models, i.e., LLaMA-2 7B and LLaMA-2 13B Touvron et al. (2023b), as well as encoder-decoder models, i.e., mT5, mT5, mT5, mT5, mT5 Xue et al. (2021). For all backbone models with <10B parameters, we conduct a full parameter fine-tuning, while for >10B parameter models (i.e., LLaMA-2 13B and mT5), we utilize a parameter efficient fine-tuning approach, LoRA Hu et al. (2022).
@@ 4.2 Multi-Phase Tuning
To develop a better instruction-tuned model, we develop the model in two phases of instruction-tuning for each backbone model. The first phase consists only of the NLP task-based prompt data with a total of 18 million instructions over-sampled from the NLP task-based prompt (§3.1) in Cendol Collection. While the second phase consists of other prompt types including general knowledge prompts (§3.2), local language generative prompts (§3.3), and human-centric prompts (§3.4) with a total of 12.8 million instructions. We divide the tuning into two phases to develop both stronger NLP task-specific and more general conversational LLMs. We denote the first phase models as Cendol-Instruct (Cendol) and the second phase models as Cendol-Chat (Cendol). We report the complete hyperparameters used in Appendix A.
@@ 6.1 Comparison with Existing LLMs
We present the results of the NLU and NLG evaluations of Cendol compared to existing LLMs in Figure 4 and Figure 5, respectively. In terms of language understanding capability, the best Cendol model (i.e., Cendol mT5) outperforms all existing LLMs both multilingual, SEA languages, and Indonesian LLMs on the comparable size by 20% weighted F1-score. Even smaller Cendol models (i.e., Cendol mT5 and Cendol mT5 with 600M and 1.2B parameters, respectively), outperform larger LLMs with 7B and 13B parameters. Similarly for language generation, we observe a huge improvement in MT, QA, and paraphrasing tasks with at least 20 increase in chrF++ Popović (2015) and SacreBLEU Post (2018), respectively. For summarization tasks, Cendol models perform similarly to Merak 7B V4 and outperform other baseline LLMs by 5% ROUGE-L. Our results signify the importance of large-scale instruction tuning to improve the zero-shot NLP capability for underrepresented regional languages.
@@ 6.2 Generalization Towards Unseen Data
| Model Type | Cendol | Cendol | Perf. |
| Cendol mT5 | 30.02 | 29.84 | -0.18 |
| Cendol mT5 | 45.08 | 35.87 | -9.21 |
| Cendol mT5 | 48.82 | 40.13 | -8.69 |
| Cendol mT5 | 58.84 | 55.79 | -3.05 |
| Cendol mT5 | 46.95 | 37.16 | -9.79 |
@@
As shown in Figure 7, the NLU performance to unseen languages follows the scaling law of LLMs. The performance improvements on NLG tasks are less apparent, moreover, no effect of scaling is observed on the translation from Indonesian to the unseen language direction. This showcases that, despite being able to better understand the unseen languages, the LLMs still have difficulty generating sentences in these unseen languages. Interestingly, we observe degradation in terms of NLU performance from the LoRA-tuned models, i.e., LLaMA2-13B and mT5, despite their increase in NLG performance.
@@
We compare the task-specific Cendol models with the general Cendol models. As shown in Table 1, the task-specific performance of Cendol models decreases significantly by up to 10% weighted F1-score. We further evaluate the models through human evaluation for both task-specific and general prompts (Figure 8). For the task-specific human evaluation, we sample 60 generation results from all the evaluated NLG tasks. For the general prompts human evaluation, we generate responses from 100 prompts that require some local knowledge in Indonesia. The responses are then rated by 3 annotators with a moderate inter-annotator agreement (=0.59). The annotation guideline is described in Appendix C.
@@
The result of human evaluation on task-specific prompts is shown in Figure 8. Cendol models significantly outperform all other models scoring a large portion of A rating compared to others. Cendol models achieve lower ratings, with LLaMA2-based Cendol performing slightly lower scores compared to the Cendol models, while mT5-based Cendol models perform on a par with other multilingual and regional LLMs.
@@
As shown in Figure 8, Cendol models fail to answer general prompts in almost all cases, while Cendol models show a trend similar to the NLP task-specific performance of Cendol where larger models perform better than a smaller model with the exception on the LoRA-tuned model, i.e., mT5. Despite the huge quality shift after the second phase of instruction-tuning, there are only several responses that are accurate and comparable to human standards (Rate A). This shows that supervised fine-tuning alone is not enough and further human-alignment tuning strategy, such as RLHF Christiano et al. (2017) or RLAIF Bai et al. (2022), is necessary to generate human-aligned responses.
@@ 6.5 Parameter Efficient Tuning Is Ineffective
We compare the effectiveness and efficiency of the parameter-efficient tuning method with LoRA Hu et al. (2022) with fully fine-tuned models with a similar training throughput. Specifically, we compare the LoRA-based Cendol mT5 model with two other models, i.e., Cendol mT5 and Cendol mT5. As shown in Table 3, the training throughput of Cendol mT5 is 1.5x higher than Cendol mT5 and is 0.34x lower than Cendol mT5. Nonetheless, in terms of other efficiency aspects, such as inference throughput and storage size, Cendol mT5 is less efficient than other models. In terms of quality, the training and evaluation losses of Cendol mT5 are much higher which leads to a worse performance downstream task performance for both NLU and NLG. These results demonstrate that despite reducing the computational resources compared to fully fine-tuned models of the same size, parameter-efficient tuning methods are less effective and less efficient compared to the smaller fully fine-tuned models. We provide the Pareto-efficiency curve of Cendol mT5 compared to the other fully fine-tuned Cendol mT5 models in Appendix E. In addition, we introduce an alternative approach for improving the modeling efficiency through vocabulary adaptation Chau et al. (2020); Poerner et al. (2020); Tai et al. (2020); Koto et al. (2021) in Appendix D.
@@ 7 Conclusion
Although our instruction-tuning data and user-oriented evaluation primarily focus on building general-purpose LLMs which are commonly expected to be able to respond interactively in a multi-turn manner. It is essential to acknowledge that our Cendol and Cendol models are not currently optimized for handling multi-turn dialogues. In other words, they are not expected to be able to engage in a continuous human-computer interaction. This can result in less coherent and less effective responses when compared to models specifically designed for a continuous human-computer interaction. Therefore, future work should focus more on developing a multi-turn dialogue system, by preserving the context and the interactions between the user and the model in previous turns and carry them to future turns.
@@ Appendix A Tuning Hyperparameters
We provide detailed hyperparameter tuning for developing Cendol and Chat for each different model architecture in Table 4.
@@ Appendix C Human Evaluation Guidelines
Appendix D Effect of Improving Token Representation
| Factor | Vocab | Vocab | Perf. |
| Model Efficiency | |||
| Token Efficiency (Ind) | 46.34 tokens | 58.87 tokens | 21.28% |
| Token Efficiency (Oth) | 52.61 tokens | 61.74 tokens | 14.79% |
| Training (per batch) | 3.14s | 3.55s | 11.50% |
| Inference (per 100 steps) | 6.63s | 8.15s | 18.71% |
| Downstream Performance | |||
| NLU Performance | 58.51 | 55.4 | 5.61% |
| NLG Performance | 45.27 | 45.79 | 1.14% |
Subword tokenization in LLMs generally produces longer sequences for low-resource language Ahia et al. (2023) which makes it less efficient. To improve the efficiency, we perform vocabulary adaptation Chau et al. (2020); Tai et al. (2020); Poerner et al. (2020) before the instruction tuning phase. Specifically, we use a subword vocabulary developed from Indonesian corpora, with the embedding initialized through averaging Koto et al. (2021), and perform the first phase instruction tuning. As shown in Table 8, the vocabulary adapted model steadily improve the token efficiency by 21.28% for Indonesian text and 14.79% for other local language text. This token efficiency results in improvements in both training and inference with around 11.50% and 18.71% efficiency improvement, respectively. Additionally, in terms of downstream performance, the vocabulary-adapted model yields a similar performance compared to the model with the original vocabulary, with improvement on the NLU tasks and % reduction on the NLG tasks. Our result suggests that vocabulary adaptation with subword averaging provides an adequately representative initialization resulting in a significantly better efficiency and similar downstream task performance after the instruction tuning phase.
Appendix E Pareto-Efficiency of Cendol Models

### v2

@@ 1 Introduction
Indonesia is the fourth most populous country in the world, with around 280 million people spread across more than 17,000 islands within a humongous area of 2 million square kilometers. With such a large archipelago surrounding the country, digital services become immensely crucial, making Indonesia the fourth largest internet user in the world, with 220 million users. Despite the huge demand, the technology supporting Indonesian digital businesses still lags compared to other much smaller countries. One aspect that it still left behind is the access to state-of-the-art large language model (LLM) technology, such as ChatGPT OpenAI (2023a) and GPT4 OpenAI (2023b). Although these LLMs support Indonesian and its local languages, these LLMs often have much weaker language representation for such low-resource and underrepresented languages Cahyawijaya et al. (2023b, a); Asai et al. (2023).
@@ 2.1 Indonesian Language Models
Various pre-trained Indonesian language models (LMs) have emerged in the past years, including IndoBERT Wilie et al. (2020); Koto et al. (2020, 2021), IndoBART Cahyawijaya et al. (2021), and IndoGPT Cahyawijaya et al. (2021). These models have smaller parameter sizes compared to recent LLMs and have primarily been evaluated only on standard NLP benchmarks. Concurrently, advancements in LLMs have led to the development of multilingual LLMs like BLOOM Scao et al. (2022) and mT5 Xue et al. (2021), which include Indonesian, Javanese, and Sundanese. Yet, they fall short of covering other underrepresented Indonesian local languages. LLaMA-2 Touvron et al. (2023b) also incorporates Indonesian, although it comprises a small portion (0.03%), diminishing its usability in the Indonesian context. Additionally, multilingual LLMs focusing on Southeast Asian languages, such as SEA-LION Singapore (2023) and SeaLLM Nguyen et al. (2023), are beginning to rise, indicating an increasing demand for refining LLMs for underrepresented languages.
@@ 3.3 Local Language Generative Prompt
To cover more underrepresented languages spoken in Indonesia, we collect Indonesian local language prompts from two sources, i.e., Wikipedia in local languages and NusaParagraph Cahyawijaya et al. (2023b). We covered 18 local languages including Sundanese (sun), Javanese (jav), Acehnese (ace), Banjarese (bjn), Buginese (bug), and Gorontalo (gor). Since many local languages in Indonesia are derived from standard Malaysian (msa), we also collect the prompt from Malaysian Wikipedia.777https://ms.wikipedia.org For the prompt from Wikipedia, we incorporate the same prompt generation strategy as in §3.2, while for the generative prompt from NusaParagraph, we invert the input and output label of the dataset to make a sentence generation task for the specified local language. In total, we collect 6.27M local language generative prompts.
@@ 4.1 Backbone Models
We prepare Cendol models of various base models to enable thorough comparison and analysis across different scales. Specifically, we train Cendol from models of different sizes, from 300M up to 13B parameters, to see the impact of size on performance. We also explore using decoder-only and encoder-only models, and lastly, we also explore using models of different origins to see if the starting base model has any impact. Finally, by producing Cendol in different configurations, users can choose their models based on needs and constraints. Specifically, we train Cendol by continuously fine-tuning decoder-only models, i.e., LLaMA-2 7B and LLaMA-2 13B Touvron et al. (2023b), as well as encoder-decoder models, i.e., mT5small, mT5base, mT5large, mT5XL, mT5XXL Xue et al. (2021). For all backbone models with <10B parameters, we conduct a full parameter fine-tuning, while for >10B parameter models (i.e., LLaMA-2 13B and mT5XXL), we utilize a parameter efficient fine-tuning approach, LoRA Hu et al. (2022).
@@ 4.2 Multi-Phase Tuning
To develop a better instruction-tuned model, we develop the model in two phases of instruction-tuning for each backbone model. The first phase consists only of the NLP task-based prompt data with a total of 18 million instructions over-sampled from the NLP task-based prompt (§3.1) in Cendol Collection. While the second phase consists of other prompt types including general knowledge prompts (§3.2), local language generative prompts (§3.3), and human-centric prompts (§3.4) with a total of 12.8 million instructions. We divide the tuning into two phases to develop both stronger NLP task-specific and more general conversational LLMs. We denote the first phase models as Cendol-Instruct (Cendolinst) and the second phase models as Cendol-Chat (Cendolchat). We report the complete hyperparameters used in Appendix A.
@@ 6.1 Comparison with Existing LLMs
We present the results of the NLU and NLG evaluations of Cendolinst compared to existing LLMs in Figure 4 and Figure 5, respectively. In terms of language understanding capability, the best Cendolinst model (i.e., Cendol mT5XL) outperforms all existing LLMs both multilingual, SEA languages, and Indonesian LLMs on the comparable size by 20% weighted F1-score. Even smaller Cendolinst models (i.e., Cendol mT5base and Cendol mT5large with 600M and 1.2B parameters, respectively), outperform larger LLMs with 7B and 13B parameters. Similarly for language generation, we observe huge improvements in MT, QA, and paraphrasing tasks with at least 20 increase in chrF++ Popović (2015) and SacreBLEU Post (2018), respectively. For summarization tasks, Cendolinst models perform similarly to Merak 7B V4 and outperform other baseline LLMs by 5% ROUGE-L. Our results signify the importance of large-scale instruction tuning to improve the zero-shot NLP capability for underrepresented regional languages.
@@ 6.2 Generalization Towards Unseen Data
| Model Type | Cendolinst | Cendolchat | Perf. |
| Cendol mT5small | 30.02 | 29.84 | -0.18 |
| Cendol mT5base | 45.08 | 35.87 | -9.21 |
| Cendol mT5large | 48.82 | 40.13 | -8.69 |
| Cendol mT5XL | 58.84 | 55.79 | -3.05 |
| Cendol mT5XXL | 46.95 | 37.16 | -9.79 |
@@
As shown in Figure 7, the NLU performance to unseen languages follows the scaling law of LLMs. The performance improvements on NLG tasks are less apparent, moreover, no effect of scaling is observed on the translation from Indonesian to the unseen language direction. This showcases that, despite being able to better understand the unseen languages, the LLMs still have difficulty generating sentences in these unseen languages. Interestingly, we observe degradation in terms of NLU performance from the LoRA-tuned models, i.e., LLaMA2-13B and mT5XXL, despite their increase in NLG performance.
@@
We compare the task-specific Cendolinst models with the general Cendolchat models. As shown in Table 1, the task-specific performance of Cendolchat models decreases significantly by up to 10% weighted F1-score. We further evaluate the models through human evaluation for both task-specific and general prompts (Figure 8). For the task-specific human evaluation, we sample 60 generation results from all the evaluated NLG tasks. For the general prompts human evaluation, we generate responses from 100 prompts that require some local knowledge about Indonesia. The responses are then rated by 3 annotators with a moderate inter-annotator agreement (=0.59). The annotation guideline is described in Appendix C.
@@
The result of human evaluation on task-specific prompts is shown in Figure 8. Cendolinst models significantly outperform all other models scoring a large portion of A rating compared to others. Cendolchat models achieve lower ratings, with LLaMA2-based Cendolchat performing slightly lower scores compared to the Cendolinst models, while mT5-based Cendolchat models perform on a par with other multilingual and regional LLMs.
@@
As shown in Figure 8, Cendolinst models fail to answer general prompts in almost all cases, while Cendolchat models show a trend similar to the NLP task-specific performance of Cendolinst where larger models perform better than a smaller model with the exception on the LoRA-tuned model, i.e., mT5XXL. Despite the huge quality shift after the second phase of instruction-tuning, there are only several responses that are accurate and comparable to human standards (Rate A). This shows that supervised fine-tuning alone is not enough and further human-alignment tuning strategy, such as RLHF Christiano et al. (2017) or RLAIF Bai et al. (2022), is necessary to generate human-aligned responses.
@@ 6.5 Parameter Efficient Tuning Is Ineffective
We compare the effectiveness and efficiency of the parameter-efficient tuning method with LoRA Hu et al. (2022) with fully fine-tuned models with a similar training throughput. Specifically, we compare the LoRA-based Cendol mT5XXL model with two other models, i.e., Cendol mT5large and Cendol mT5XL. As shown in Table 3, the training throughput of Cendol mT5XXL is 1.5x higher than Cendol mT5XL and is 0.34x lower than Cendol mT5large. Nonetheless, in terms of other efficiency aspects, such as inference throughput and storage size, Cendol mT5XXL is less efficient than other models. In terms of quality, the training and evaluation losses of Cendol mT5XXL are much higher which leads to a worse performance downstream task performance for both NLU and NLG. These results demonstrate that despite reducing the computational resources compared to fully fine-tuned models of the same size, parameter-efficient tuning methods are less effective and less efficient compared to the smaller fully fine-tuned models in the case of language adaptation. We provide the Pareto-efficiency curve of Cendol mT5XXL compared to the other fully fine-tuned Cendol mT5 models in Appendix D. As an alternative solution for language adaptation, we introduce an alternative approach for improving the modeling efficiency through vocabulary adaptation Chau et al. (2020); Poerner et al. (2020); Tai et al. (2020); Koto et al. (2021).
@@
| Factor | Vocabind | Vocaborig | Perf. |
| Model Efficiency | |||
| Token Efficiency (Ind) | 46.34 tokens | 58.87 tokens | 21.28% |
| Token Efficiency (Oth) | 52.61 tokens | 61.74 tokens | 14.79% |
| Training (per batch) | 3.14s | 3.55s | 11.50% |
| Inference (per 100 steps) | 6.63s | 8.15s | 18.71% |
| Downstream Performance | |||
| NLU Performance | 58.51 | 55.4 | 5.61% |
| NLG Performance | 45.27 | 45.79 | 1.14% |
Subword tokenization in LLMs generally produces longer sequences for low-resource language Ahia et al. (2023) which makes it less efficient. By performing vocabulary adaptation Chau et al. (2020); Tai et al. (2020); Poerner et al. (2020), we can improve the efficiency during the instruction tuning phase. Specifically, we use a subword vocabulary developed from Indonesian corpora, with the embedding initialized through averaging Koto et al. (2021), and perform the first phase instruction tuning. As shown in Table 4, the vocabulary adapted model steadily improve the token efficiency by 21.28% for Indonesian text and 14.79% for other local language text. This token efficiency results in improvements in both training and inference with around 11.50% and 18.71% efficiency improvement, respectively. Additionally, in terms of downstream performance, the vocabulary-adapted model yields a similar performance compared to the model with the original vocabulary, with improvement on the NLU tasks and % reduction on the NLG tasks. Our result suggests that vocabulary adaptation with subword averaging provides an adequately representative initialization resulting in a significantly better efficiency and similar downstream task performance after the instruction tuning phase.
@@ 7 Conclusion
Acknowledgements
We thank Direktorat Jenderal Pendidikan Tinggi, Riset, dan Teknologi Kementerian Pendidikan, Kebudayaan, Riset, dan Teknologi (Ditjen DIKTI), Indonesia for partially supporting the computing resources used in this project. This work is also supported by the National Research Foundation, Singapore under its AI Singapore Programme; PhD Fellowship Award, the Hong Kong University of Science and Technology; and PF20-43679 Hong Kong PhD Fellowship Scheme, Research Grant Council, Hong Kong.
@@ Acknowledgements
Although our instruction-tuning data and user-oriented evaluation primarily focus on building general-purpose LLMs which are commonly expected to be able to respond interactively in a multi-turn manner. It is essential to acknowledge that our Cendolinst and Cendolchat models are not currently optimized for handling multi-turn dialogues. In other words, they are not expected to be able to engage in a continuous human-computer interaction. This can result in less coherent and less effective responses when compared to models specifically designed for a continuous human-computer interaction. Therefore, future work should focus more on developing a multi-turn dialogue system, by preserving the context and the interactions between the user and the model in previous turns and carry them to future turns.
@@ Appendix A Tuning Hyperparameters
We provide detailed hyperparameter tuning for developing Cendolinst and Chat for each different model architecture in Table 5.
@@ Appendix C Human Evaluation Guidelines
Appendix D Pareto-Efficiency of Cendol Models